Problem 4 Erics preferences for goods x and y are represented by the following utility function: U(X,Y) = 4X +5Y. The pricecan someone explain this step by step? especially don't understand how they got m/px? why we can't use the lagrange. and not sure how they drew the graph for this question. SOMEONE PLEASE HELP MIDTERM SOON AND WILL GIVE BIG THUMBS UP!!!! confused where 4x20+5x0 is coming from
Answer #1

a)

U(X, Y) = 4X + 5Y

X and Y are perfect substitutes of each other. There exists a corner solution in this case. Eric either spends his entire income on good X or on good Y.

Marginal Rate of Substitution = MUx / MUy

MUx = Derivative of U wrt X = 4

MUy = Derivative of U wrt Y = 5

Hence,  Marginal Rate of Substitution = MUx / MUy = 4/5

Since, it is a constant numbere, so MRS in constant in X

b.

Constrained Optimization problem for Eric.

Maximize U(X, Y) subject to Budget Constraint.

Let Budget Constraint: M = Px*X + Py*Y

Px = 2, Py = 3 , M = 10

Budget Constraint: 10 = 2X + 3Y

Hence, the Constrained Optimization problem is following:

Maximize U(X, Y) st 10 >= 2X + 3Y

c.

Now, the optimal bundle will have 0 units of either of the two goods are both are perfect substitute of each other. A slight fall in the price of good X would induce Eric to demand all X and no Y.

Similarly, A slight fall in the price of good Y would induce Eric to demand only Y and 0 units of X

So, the optimal bundle is either (X* , 0) or (0, Y*)

This depends on the MRS and Price Ratio whichever is greater.

Let say MRS > Price Ratio

MUx / MUy > Px / Py

MUx / Px > MUy / Py

Eric derives more utility from good X and therefore will consume only X

Similarly,

Let say MRS < Price Ratio

MUx / MUy < Px / Py

MUx / Px < MUy / Py

Eric derives more utility from good Y and therefore will consume only Y

In our case, Price Ratio = 2/3 = .67

And MRS = 4/5 = 0.8

Since, MRS > Price Ratio. Eric will consume only X and 0 units of Y

So, Y = 0. Put this Y in Budget Constraint: 10 = 2X + 3Y

10 = 2X + 3*0

X* = 5

So, OPTIMAL BUNDLE: (X*, Y*) = (5 , 0)

d)

Eric's utility at this optimal bundle:

U = 4X + 5Y

U = 4*5 + 5*0

U = 20

Now, Eric's budget income = 20

Since, all the Eric's income has been exhausted in purchasing X and Y, there will no other bundle that would be feasible for Eric.

e)

Now, Budget Line: 10 = 2X + 3Y

3Y = 10 - 2X

Y = 10/3 - 2/3X

Hence, the slope of budget line = 2/3

Here, the Y intercept = 10/3 ( X = 0)

and X intercept = 5 (Put Y =0)

Now, U = 4X + 5Y

5Y = U - 4X

Y = 1/5*U - 4/5X

Slope of IC = 4/5 = 0.8

Slope of IC > Slope of budget line

Therefore, the ICs will be steeper than the budget line

Now, at optimal bundle Y =0

and X* = 5

10/3 T → IC (ii) Budget Line (i) + ii) Optimal Bundle

If you are unclear on something, then please ask in comments.

If you liked the answer, then please upvote. Would be motivating for me. Thanks